# A Generic Machine Learning Framework for Radio Frequency Fingerprinting

### v1

@@ 5.4.6 Use Case 6: OSR-EDA for DMR data
In this section, we present experimental results for the OSR-EDA task using the DMR dataset. Any EDA model can compare input pairs without their emitter labels, meaning we can train an EDA model on known emitters and evaluate the model on potentially unseen emitters. This is similar to the methodology in Section 4.4 but for EDA. Therefore, we trained an EDA model on the DMR dataset using emitter data from six handsets, leaving the remaining two emitters’ data held out. We then evaluated the held out data using the trained EDA model to evaluate its OS capability. In addition, we adjusted a proportion parameter $p$ of the training data to understand how much data would be needed to obtain sufficiently good performance (i.e. $p=1$ uses all of training dataset for training, $p=0.5$ uses half of the training dataset, etc) on the held out emitter data. Dropping data from the training dataset was done at random.
@@
Figures 11, 12 and 13 shows a summary of the evaluation of emitter held out data using a trained EDA parameterized with a BCNN architecture for $p=0.5,0.25,0.025$ respectively. Notice how the t-SNE embeddings of each emitter are distinctly separate and the distance distributions have some overlap between them but overall are quite well separated. The silhouette score is also high for cluster number 2, giving more indication on the structure of the evaluated held out emitter data. As $p$ is reduced, however, the performance deterioriates but is minor. These are promising experimental results towards general purpose RFF models.
6 Conclusions
@@
This paper introduces a framework for building a general-purpose machine learning RFF model. Within this, we describe a variety of RF fingerprint-related downstream tasks. We then demonstrate the four tasks SEI, EDA, RFEC and OSR-EDA individually using three real-world RF datasets (i.e. DMR for SIGINT, AIS for maritime surveillance and drone uplink/downlink signals for C-UAS). From this, we observe that low complexity models deliver a competitive performance when compared with their more complex counterparts across all tasks and datasets. This supports the implementation of high performance data-driven RFF techniques on edge devices, which are typically characterised by a low size, weight, power and cost hardware profile. The presented generic ML-based framework for RF Fingerprinting, single task formulations, and experimental work using real RF data can stimulate future research in this domain.
@@ Appendix A Data Transformation Operator for EDA
Here we introduce a sampling algorithm $\mathcal{M}$ for the EDA task. We assume that $\mathcal{D}^{(a)}=\mathcal{M}(\mathcal{D}^{(i)})=\{({\bf{x}}_{j}^{(a),1},{\bf{x}}_{j}^{(a),2},y^{(a)}_{j})\}_{j=1}^{n_{a}}$ where we choose $n_{a}=\gamma\ll\frac{n_{i}(n_{i}-1)}{2}$, $\gamma>0$. Setting this condition on the number of samples in the EDA dataset is to avoid the combinatorial explosion of pairs when $n_{i}$ is large. This way, we can sample from all possible pairs without completion, which could lead to a large computational overhead for some problems. For EDA labelling, we say that:
@@
In general, the EDA problem has class imbalance because as the number of classes $\mathcal{V}$ increases, the ratio of matched to unmatched pairs declines. The number of matched $n_{matched}=\mathcal{V}$ and the number of unmatched $n_{unmatched}=\frac{\mathcal{V}(\mathcal{V}-1)}{2}$. This allows us to define the ratio $n_{r}=\frac{n_{matched}}{n_{unmatched}}=\frac{\mathcal{V}}{\frac{1}{2}\mathcal{V}(\mathcal{V}-1)}=\frac{2}{\mathcal{V}-1}$. This can inform us on how we should structure the sampling, in terms of proportionality.
We now define a function:
which determines the number of input pairs for every class pairing $(i,j)$, $i,j\in\{1,\cdots,\mathcal{V}\}$, that is to be included in ${\mathcal{D}}^{(a)}$. The parameter $\alpha\in(0,1]$ is a ratio of matched to unmatched pairs and $\gamma$ is the total number of samples that is sampled from $\mathcal{D}^{(i)}$. We can structure $\mathcal{D}^{(i)}$ so that
where $n_{i}^{k}$ is the number of samples in dataset $\mathcal{D}^{(i)}_{k}$. Algorithm 2 shows a recipe for computing the dataset $\mathcal{D}^{(a)}$ using $\mathcal{D}^{(i)}$.

### v2

@@ VI-F Use Case 6: OS-EDA for DMR data
We present here open-set EDA task results with the DMR dataset. An EDA model can compare input pairs without their emitter labels, meaning we can train an EDA model on known emitters and evaluate it on potentially unseen emitters, hence OS-EDA. This is similar to the methodology in Section IV-D, however for EDA. We trained an EDA model on the DMR dataset using emitter data from six handsets. We held out the data of the remaining two emitters’ to be used for testing to assess the model’s OS capability. In addition, we adjusted a proportion parameter $p$ of the training data to understand how much data would be needed to obtain sufficiently good performance (i.e. $p=1$ uses all of training dataset, $p=0.5$ uses half of it, etc) on the two held out emitters’ data.
@@
Figures 11, 12 and 13 display a summary of the evaluation on data for held out emitters with a BCNN architecture for EDA with $p=0.5,0.25,0.025$ respectively. It can be noted from all the three figures that the t-SNE embeddings of each emitter are distinguishable. Whilst the distance distributions have some overlap between them, overall they are well separated. The silhouette score is also high for cluster number 2, giving more indication on the structure of the evaluated (two) emitters’ data. As $p$ is reduced, the performance marginally deteriorates. These are promising experimental results on training generic RFF models.
VII Conclusions
@@
This paper introduces a framework for building a general-purpose machine learning RF fingerprinting model. Within this, we describe a variety of RF fingerprint-dependent downstream tasks and demonstrate the performance of a number of them, namely SEI, EDA, RFEC and OS-EDA, using three real RF datasets (i.e. DMR for SIGINT, AIS for spaceborne surveillance and drone RF signals for C-UAS). From this, we observe that low complexity ML model architectures deliver a competitive performance when compared with their more complex counterparts across all tasks and datasets. This supports the implementation of high performance data-driven RFF techniques on edge devices, which are typically characterised by a low size, weight, power and cost hardware profile. The presented generic ML-based framework for RF fingerprinting and presented results can serve as an impetus to future research on data-driven RFF.
@@ Appendix A Data Transformation Operator for EDA
We introduce in this Appendix a sampling algorithm $\mathcal{M}$ for the EDA task. We assume that $\mathcal{D}^{(a)}=\mathcal{M}(\mathcal{D}^{(i)})=\{({\bf{x}}_{j}^{(a),1},{\bf{x}}_{j}^{(a),2},y^{(a)}_{j})\}_{j=1}^{n_{a}}$ where we choose $n_{a}=\gamma\ll\frac{n_{i}(n_{i}-1)}{2}$, $\gamma>0$. Setting this condition on the number of samples in the EDA dataset is to avoid the combinatorial explosion of pairs when $n_{i}$ is large. This way, we can sample from all possible pairs without completion, which could lead to a large computational overhead for some problems. For EDA labeling, we have:
@@
In general, the EDA problem has class imbalance because as the number of classes $\mathcal{V}$ increases, the ratio of matched to unmatched pairs declines. The number of matched $n_{matched}=\mathcal{V}$ and the number of unmatched $n_{unmatched}=\frac{\mathcal{V}(\mathcal{V}-1)}{2}$. This allows us to define the ratio $n_{r}=\frac{n_{matched}}{n_{unmatched}}=\frac{\mathcal{V}}{\frac{1}{2}\mathcal{V}(\mathcal{V}-1)}=\frac{2}{\mathcal{V}-1}$. This can inform how to structure the sampling, in terms of proportionality. We now define a function:
which determines the number of input pairs for every class pairing $(i,j)$, $i,j\in\{1,\cdots,\mathcal{V}\}$, that is to be included in ${\mathcal{D}}^{(a)}$. The parameter $\alpha\in(0,1]$ is a ratio of matched to unmatched pairs and $\gamma$ is the total number of samples that is sampled from $\mathcal{D}^{(i)}$. We can structure $\mathcal{D}^{(i)}$ such that
where $n_{i}^{k}$ is the number of samples in $\mathcal{D}^{(i)}_{k}$. Algorithm 2 shows a recipe for computing dataset $\mathcal{D}^{(a)}$ using $\mathcal{D}^{(i)}$.